# Robotic Sorting Systems: Robot Management and Layout Design Optimization

### v1

@@ 8 Numerical Examples
where $M_{i}$ denotes the unit cost of component $i$ in a month. We estimate the monthly warehousing rental cost using JD Logistics’ financial reports for the first three quarters of 2023. Labor cost is sourced from data related to warehousing job recruitment in Beijing, obtained from website for job posts (58.com). The data regarding the equipment cost of loading stations and the rental cost of robots are obtained from Geekplus company. The ratio of peak month $\theta$ is set to 1/6, which means that every quarter has a two-week-long promotion campaign in average. Meanwhile, we set ${T}^{L}=0.8\cdot{T}^{H}$ to ensure that there is no enormous space waste during the off-peak period. The default parameter of experiments is shown in table 5 and the configurations of RC-S keeps the same with table 2. Our algorithms are coded in Python 3.11 with the solver Gurobi Optimizer 9.0, and all experiments are conducted on a personal computer running Windows 11 with Intel i7-9700F CPU and 16GB RAM. All processes are run single-threaded.
Table 5: Default parameter in sensitivity analysis
@@ 8.1 Site rental cost
We begin by investigating the impact of site rental cost under two different scenarios: one with $N_{o}=100$ (scenario 1) and another with $N_{o}=400$ (scenario 2). The monthly rental cost of site floor space per square meter varies from $10\ CHY$ to $30\ CHY$. The results are presented in Figure 16(a) and Figure 16(b). The findings reveal that:
@@
  • Comparing the data for $T^{H}=15,000$ and $T^{H}=30,000$ in Table 6, the total costs for the latter are nearly three times those for the former. This indicates that due to the simultaneous growth of $C_{s}$ and $C_{o}$, the total costs growth rate of the RSS system exceeds the throughput growth rate.
@@
8.2 Labor cost
In a similar manner as our previous experiments, we now explore the influence of different labor costs on the total costs in two scenarios with $N_{o}=100$ and $N_{o}=400$. The monthly labor cost of each worker varies from $5,000\ CHY$ to $12,000\ CHY$. The proportions of labor costs and robot rental costs are both visualized in Figure 17(a) and Figure 17(b). Our observations are concluded as follows:
@@ 8.2 Labor cost
Additionally, the cost structure undergoes notably changes with varying sorting demands. In scenarios with high sorting demand, the proportion and growth rate of robot rental cost become more significant. This reveals a distinctive feature of RSS systems: Robots, as carriers of parcels, become less efficient in larger sites because they have to cover increased distances. Merely enhancing management methods is insufficient to tackle the challenge of diminishing traffic efficiency. Today’s commonly used conveyor sorters, equipped with high-power motors, necessitate substantial initial investments but experience only slight increments in total costs as sorting demands grow (Zou et al.,, 2021). Furthermore, the optimal system configurations of a conveyor sorter are not easily affected by fluctuations in the unit price of cost components (Russell and Meller, 2003b, ), which makes it suitable for long-term operation in large-scale sorting scenarios. Therefore, we conclude the following insight regarding the application of RSS:
@@ 9 Conclusion
In the operations stage, efficient robot scheduling emerges as a crucial element for unlocking the business value of RSS. We propose an innovate RC-S scheme, which serves as an framework for managing a considerable number of robots simultaneously and efficiently. We provide a detailed description of the composition of phases and the method for setting parameters in RC-S. Additionally, we present a mathematical programming model FPA to minimize the travel costs at each cycle, and a heuristic algorithm as online solver. Subsequently, a theoretical analysis of the efficiency of RC-S is conducted, exploring the impact of various system configurations on throughput. In the validation section, we first compare our proposed traffic management framework with other common framework based on A*. The simulation results indicate that our control method achieves a high level of performance and computational efficiency. Furthermore, we validate the throughput estimation formula. Results show that the analytical model fits the real traffic of robots well.
In the site planning stage, warehouse managers need to make decisions that involve a trade-off between the initial site investment $C_{s}$, and the discounted expected operations costs $C_{o}$. We analyze the cost composition of the RSS, and propose a layout design optimization model LDP that minimizes the total system costs. Through a detailed sensitivity analysis examining the cost dynamics of site rental and labor, our findings demonstrate a distinct cost distribution based on throughput levels. Specifically, at lower throughput, the rental cost for the site represents a considerable proportion of the total costs. However, at higher throughput, the expenditure on labor emerges as the predominant cost element. Results also validate the discussions on the properties of our model. The key insights can guide managers in understanding the investment and returns of applying the RSS, thereby reducing the total costs of the warehouse.
@@
$\mathcal{G}$
graph constituted of all nodes and links
$\mathcal{V}_{c}$
set of all conflict nodes
$\mathcal{V}_{u}$
set of all unloading nodes
$\mathcal{A}$
set of all links
$\mathcal{S}$
set of all loading stations
$\mathcal{O}$
set of all outlets
$\mathcal{E}$
set of all cycles
$\mathcal{R}$
set of all feasible path
$\mathcal{R}(i,j,k)$
set of feasible path connecting loading station $i$ and $j$, passing outlet $k$
$\mathcal{R}_{i},k$
set of feasible path starting at loading station $i$ and passing by outlet $k$
^$\hat{\mathcal{R}}_{i}$
set of feasible path ending at loading station $i$
$\mathcal{I}_{x}$
set of locations of loading stations on the top and bottom sides
$\mathcal{I}_{y}$
set of locations of loading stations on the left and right sides
@@
$D$
length of each grid in the sorting zone
$W_{w}$
width of waiting zone
$W_{l}$
width of loading zone
$C_{f}$
the facility costs of RSS
$C_{o}$
the operations cost of RSS
$C_{d}$
the total costs of RSS
$P_{f}$
cost of facility per square floor space
$P_{l}$
cost of equipment in one loading station
$P_{w}$
cost of labor per person per month
$P_{r}$
cost of robot per vehicle per month
${T}_{H},{T}_{L}$
target throughput level in peak/off-peak season
$\theta$
Ratio of peak seasons
$N_{o}$
minimum number of outlets required to meet the sorting category demands
$\gamma$
weight of operations costs in optimization model
$\tau_{c}$
time interval of VPs in RC-S
$\tau_{e}$
travel time of VPs on each link in RC-S
$v_{VP}$
speed of VPs
$v_{max}$
maximum speed of robots
$c_{max}$
maximum acceleration/deceleration of robots
$r_{ls}$
maximum loading rate of a loading station
$c_{i,k}^{r}$
travel time of feasible path $r\in R_{i,k}$
$\hat{c}_{i,k}$
penalty of delay of a sorting task from loading station $i$ to outlet $k$
$\delta_{i,k}^{r,\nu,e}$
incidence between feasible path $r\in R_{i,k}$, node $\nu$ and cycle $e$
$N_{\nu}^{e}$
remaining capacity in node $\nu$ in cycle $e$
$d_{i,k}$
sorting demand from loading station $i$ to outlet $k$
@@
$x_{i,k}^{r}$
decision on whether feasible path $r\in R_{i,k}$ is reserved in current cycle
$\hat{x}_{i,k}$
decision on whether the robot located at loading station $i$ should wait until the next cycle, with the target outlet $k$
$n_{h},n_{v}$
number of horizontal and vertical aisles in sorting zone
$n_{l}$
number of loading stations in loading zone
$n_{w}^{H},n_{w}^{L}$
number of workers in peak/off-peak season
$n_{r}^{H},n_{r}^{L}$
number of robots in peak/off-peak season
@@
B.1 Derivation of Maximum Travel Distance in Constraint 5
@@
B.2 Derivation of Average Travel Distance in Section 5
@@ B.4 Proof of Proposition 1
$\tilde{T}_{M}(n_{h},n_{v},n_{l})=\frac{D}{\tau_{e}}\cdot\frac{n_{h}(n_{v}-1)+n_{v}(n_{h}-1)}{2[a+b(n_{h}+n_{v})]}\cdot\frac{\frac{2n_{l}}{n_{h}+n_{v}}-{(\frac{n_{l}}{n_{h}+n_{v}})}^{2}}{\bar{l}(n_{h},n_{v},\frac{n_{l}}{n_{h}+n_{v}})}$
@@
$\frac{\partial\tilde{T}_{M}(n_{h},n_{v},n_{l})}{\partial}=\frac{D}{\tau_{e}}\cdot\frac{n_{h}(n_{v}-1)+n_{v}(n_{h}-1)}{2[a+b(n_{h}+n_{v})]}\cdot\frac{1}{n_{h}+n_{v}}\cdot\frac{\partial G_{n_{h},n_{v}}(\alpha)}{\partial\alpha}$
@@
$s.t.\quad$ $y_{i}\geq g_{i}(\omega_{k})+\nabla{g_{i}(\omega_{k})}^{T}d,\quad i=1,2,\cdots,6$ (C-1)
@@
$y_{i}\geq 0,\quad i=1,2,\cdots,6$ (C-3)
where $\omega=\{n_{h},n_{v},n_{w}^{H},n_{r}^{H},n_{w}^{L},n_{r}^{L}\}$ is the set of all decision variables. $\Delta_{k}$ is the bound of $d$ in iteration $k$. If we impose the constraint that the number of robots is less than the available VPs (considering the parking demand of robots during downtime, this setting is reasonable), constraints in layout design optimization model LDP are replaced with six inequalities in the LP-S:
$g_{1}(\omega)=n_{r}^{H}-\beta(n_{h},n_{v})\cdot(1-{(1-\frac{n_{w}^{H}}{n_{h}+n_{v}})}^{2})\cdot\frac{2\tau_{e}\cdot[n_{h}(n_{v}-1)+n_{v}(n_{h}-1)]}{\tau_{c}}\leq 0$ (C-4)
$g_{2}(\omega)=n_{r}^{L}-\beta(n_{h},n_{v})\cdot(1-{(1-\frac{n_{w}^{L}}{n_{h}+n_{v}})}^{2})\cdot\frac{2\tau_{e}\cdot[n_{h}(n_{v}-1)+n_{v}(n_{h}-1)]}{\tau_{c}}\leq 0$ (C-5)
$g_{3}(\omega)=-\frac{D\cdot n_{r}^{H}}{\tau_{e}\cdot E[l(n_{h},n_{v},\frac{n_{w}^{H}}{n_{h}+n_{v}})]}+{T}^{H}\leq 0$ (C-6)
$g_{4}(\omega)=-\frac{D\cdot n_{r}^{L}}{\tau_{e}\cdot E[l(n_{h},n_{v},\frac{n_{w}^{L}}{n_{h}+n_{v}})]}+{T}^{L}\leq 0$ (C-7)
$g_{5}(\omega)=-(n_{h}-1)(n_{v}-1)+N_{o}\leq 0$ (C-8)
$g_{6}(\omega)=n_{w}^{H}-n_{h}-n_{v}\leq 0$ (C-9)
To avoid waste of resources, the number of loading station is always equal to the peak number of workers, thus $n_{l}$ is replaced by $n_{w}^{H}$ in the constraints. The details of solving PLSP is shown in Algorithm C-1. It could be roughly divided into 2 steps in each iteration: (1) Generate LP-S and obtain the step size $d_{k}$, determine whether to stop; (2) Adjust step bounds for the next iteration.
@@
$C_{E}(\omega_{k})=C_{d}(\omega_{k})+\mu\cdot\sum_{i=1}^{6}max\{0,g_{i}(\omega_{k})\}$ (C-10)
$C_{EL_{k}}(\omega_{k})=C_{d}(\omega_{k})+\nabla{C_{d}(\omega_{k})}^{T}\cdot d_{k}+\mu\cdot\sum_{i=1}^{6}max\{0,g_{i}(\omega_{k})+{\nabla g_{i}(\omega_{k})}^{T}\cdot d_{k}\}$ (C-11)
$R_{k}=\frac{C_{E}(\omega_{k})-C_{E}(\omega_{k}+d_{k})}{C_{EL_{k}}(\omega_{k})-C_{EL_{k}}(\omega_{k}+d_{k})}$ (C-12)
where $\mu$ is a large enough constant. Given that the variables in the original problem are integers, we introduced a rounding-up step for $\omega_{k}$ and the bounds in the search process. Additionally, due to the relatively limited solution space, the number of iterations in experiments remained below 50, and the total solving time was in the order of seconds. Therefore, specific details are not explicitly described in the main text.

### v2

@@ 8 Numerical Examples
where $M_{i}$ denotes the unit cost of component $i$ in a month. We estimate the monthly warehousing rental cost using JD Logistics’ financial reports for the first three quarters of 2023. Labor cost is sourced from data related to warehousing job recruitment in Beijing, obtained from website for job posts (58.com). The data regarding the equipment cost of loading stations and the rental cost of robots are obtained from Geekplus company.
We assume that the warehouse only expands its sorting capacity during a few major shopping events, meaning there are two typical levels of sorting demand, namely the average off-peak season demand ${T}^{L}$ and the average peak season demand ${T}^{H}$, with ${T}^{L}=0.8\cdot{T}^{H}$. The ratios of the two periods are set as $(\theta^{L},\theta^{H})=(5/6,1/6)$, which means that each quarter typically includes a two-week-long shopping event. The default values of parameters is shown in table 6 and the configurations of RC-S keeps the same with table 2. Our algorithms are coded in Python 3.11 with the solver Gurobi Optimizer 9.0, and all experiments are conducted on a personal computer running Windows 11 with Intel i7-9700F CPU and 16GB RAM. All processes are run single-threaded.
Table 6: Default values of parameters in section 8
@@ 8.1 Site rental cost
We begin by investigating the impact of site rental cost under two different scenarios: one with $N_{o}=100$ (scenario 1) and another with $N_{o}=400$ (scenario 2). The monthly rental cost of site floor space per square meter varies from $10\ CNY$ to $30\ CNY$. The results are presented in Figure 16(a) and Figure 16(b). The findings reveal that:
@@
  • Comparing the data for $T^{H}=15,000$ and $T^{H}=30,000$ in Table 7, the total costs for the latter are nearly three times those for the former. This indicates that due to the simultaneous growth of $C_{f}$ and $C_{o}$, the total costs growth rate of the RSS system exceeds the throughput growth rate.
@@
8.2 Labor cost
In a similar manner as our previous experiments, we now explore the influence of different labor costs on the total costs in two scenarios with $N_{o}=100$ and $N_{o}=400$. The monthly labor cost of each worker varies from $5,000\ CNY$ to $12,000\ CNY$. The proportions of labor costs and robot rental costs are both visualized in Figure 17(a) and Figure 17(b). Our observations are concluded as follows:
@@ 8.2 Labor cost
Additionally, the cost structure undergoes notably changes with varying sorting demands. In scenarios with high sorting demand, the proportion and growth rate of robot rental cost become more significant. This reveals a distinctive feature of RSS systems: Robots, as carriers of parcels, become less efficient in larger sites because they have to cover increased distances. Merely enhancing management methods is insufficient to tackle the challenge of diminishing traffic efficiency. Today’s commonly used conveyor sorters, equipped with high-power motors, necessitate substantial initial investments but experience only slight increments in total costs as sorting demands grow (Zou et al.,, 2021). Furthermore, the optimal system configurations of a conveyor sorter are not easily affected by fluctuations in the unit price of cost components (Russell and Meller,, 2003), which makes it suitable for long-term operation in large-scale sorting scenarios. Therefore, we conclude the following insight regarding the application of RSS:
@@ 9 Conclusion
In the operations stage, efficient robot scheduling emerges as a crucial element for unlocking the business value of RSS. We propose an innovative RC-S scheme, which serves as a framework for managing a considerable number of robots simultaneously and efficiently. We provide a detailed description of the composition of phases and the method for setting parameters in RC-S. Additionally, we present a mathematical programming model FPA to minimize the travel costs at each cycle, and a heuristic algorithm as online solver. Subsequently, a theoretical analysis of the efficiency of RC-S is conducted, exploring the impact of various system configurations on throughput. In the validation section, we first compare our proposed traffic management framework with classical cooperative A* algorithm. The simulation results indicate that our control method achieves a higher level of performance and computational efficiency. Furthermore, we validate the throughput estimation formula. Results show that queueing models can experience significant distortion in some scenarios, while our model fits the real traffic flows of robots more precisely.
In the site planning stage, warehouse managers need to make decisions that involve a trade-off between the initial investment of facility $C_{f}$, and the discounted operations costs $C_{o}$. We analyze the cost composition of the RSS, and propose a layout design optimization model LDP that minimizes the total system costs. Through a detailed sensitivity analysis examining the cost dynamics of site rental and labor, we investigate cost proportions under distinct throughput levels. Specifically, at lower throughput, the rental cost for the site represents a considerable proportion of the total costs. However, at higher throughput, the expenditure on labor emerges as the predominant cost element. Results also validate the discussions on the properties of our model. The key insights can guide managers in understanding the investment and returns of applying the RSS, thereby reducing the total costs of the warehouse.
@@
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used ChatGPT in order to check for grammatical errors and improve readability. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
@@
$\mathcal{G}$ graph constituted of all nodes and links
$\mathcal{V}_{c}$ set of conflict nodes
$\mathcal{V}_{u}$ set of unloading nodes
$\mathcal{V}_{e}$ set of entrance/exit nodes
$\mathcal{E}$ set of all edges
$\mathcal{L}$ set of all loading stations
$\mathcal{O}$ set of all outlets
$\mathcal{C}$ set of all cycles
$\mathcal{R}$ set of all feasible path
$\mathcal{R}(i,j,k)$ set of feasible path connecting loading station $i$ and $j$, passing outlet $k$
$\mathcal{R}_{i},k$ set of feasible path starting at loading station $i$ and passing by outlet $k$
^$\hat{\mathcal{R}}_{i}$ set of feasible path ending at loading station $i$
$\mathcal{I}_{x}$ set of locations of loading stations on the top and bottom sides
$\mathcal{I}_{y}$ set of locations of loading stations on the left and right sides
$\mathcal{S}$ set of operating periods
@@
$D$ length of each grid in the sorting zone
$W_{w}$ width of waiting zone
$W_{l}$ width of loading zone
$C_{f}$ the facility costs of RSS
$C_{o}$ the operations cost of RSS
$C_{d}$ the total costs of RSS
$P_{f}$ cost of facility per square floor space
$P_{l}$ cost of equipment in one loading station
$P_{w}$ cost of labor per person per month
$P_{r}$ cost of robot per vehicle per month
${T}^{\sigma}$ target throughput level in operating period $\sigma$
$\theta^{\sigma}$ Ratio of operating period $\sigma$
$N_{o}$ minimum number of outlets required to meet the sorting category demands
$\gamma$ weight of operations costs in optimization model
$\tau_{c}$ time interval of VPs in RC-S
$\tau_{e}$ travel time of VPs on each link in RC-S
$v_{VP}$ speed of VPs
$v_{max}$ maximum speed of robots
$c_{max}$ maximum acceleration/deceleration of robots
$r_{l}$ maximum loading rate of a loading station
$c_{i,k}^{r}$ travel time of feasible path $r\in R_{i,k}$
$\hat{c}_{i,k}$ penalty of delay of a sorting task from loading station $i$ to outlet $k$
$\delta_{i,k}^{r,\nu,l}$ incidence between feasible path $r\in R_{i,k}$, node $\nu$ and cycle $l$
$N_{\nu}^{l}$ remaining capacity in node $\nu$ in cycle $l$
$d_{i,k}$ sorting demand from loading station $i$ to outlet $k$
@@
$x_{i,k}^{r}$ decision on whether feasible path $r\in R_{i,k}$ is reserved in current cycle
$\hat{x}_{i,k}$ decision on whether the robot located at loading station $i$ should wait until the next cycle, with the target outlet $k$
$n_{h},n_{v}$ number of horizontal and vertical aisles in sorting zone
$n_{l}$ number of loading stations in loading zone
$n_{w}^{\sigma}$ number of workers in operating period $\sigma$
$n_{r}^{\sigma}$ number of robots in operating period $\sigma$
@@
B.1 Proof of Proposition 1
(i):
By definition 1, a feasible path of RC-S connects two active loading stations. If a feasible path has no more than two turns, all its turning points are within the blue region; otherwise, its endpoint will fall into a non-active loading station. The movement trajectories of occupied VPs will not cross the boundaries of the blue region; therefore, their upper limit is $\kappa$. Condition (i) is sufficient.
(ii):
We derive the average road length that each loading station can be allocated, noted as $\tilde{d}_{ls}$:
$\tilde{d}_{ls}(\alpha)$ $=$ $\frac{\big{(}1-{(1-\alpha)}^{2}\big{)}\cdot n_{h}n_{v}}{\frac{\alpha(n_{h}+n_{v})}{2}}$
$=$ $\frac{(2-\alpha)\cdot 2n_{h}n_{v}}{n_{h}+n_{v}}\geq(2-\alpha)\cdot min\{n_{h},n_{v}\}$
When the average travel distance of a task is less than $\tilde{d}_{ls}$, the proportion of the union of segments passed by the occupied VPs in the total set of segments is less than $\kappa$. Condition (ii) is sufficient.
$\square$
B.2 Derivation of Maximum Travel Distance in Constraint 5
@@
B.3 Derivation of Average Travel Distance in Section 5
@@ B.5 Proof of Proposition 2
$\tilde{T}_{M}(n_{h},n_{v},n_{l})=\frac{D}{\tau_{e}}\cdot\frac{n_{h}n_{v}}{a+b(n_{h}+n_{v})}\cdot\frac{\frac{2n_{l}}{n_{h}+n_{v}}-{(\frac{n_{l}}{n_{h}+n_{v}})}^{2}}{\bar{l}(n_{h},n_{v},\frac{n_{l}}{n_{h}+n_{v}})}$
@@
$\frac{\partial\tilde{T}_{M}(n_{h},n_{v},n_{l})}{\partial n_{l}}=\frac{D}{\tau_{e}}\cdot\frac{n_{h}n_{v}}{a+b(n_{h}+n_{v})}\cdot\frac{1}{n_{h}+n_{v}}\cdot\frac{\partial G_{n_{h},n_{v}}(\alpha)}{\partial\alpha}$
@@
$s.t.\quad$ $y_{i}\geq g_{i}(\omega_{k})+\nabla{g_{i}(\omega_{k})}^{T}d,\quad i=1,2,\cdots,2\left|S\right|+2$ (C-1)
@@
$y_{i}\geq 0,\quad i=1,2,\cdots,2\left|S\right|+2$ (C-3)
where $\omega=\{n_{h},n_{v},n_{w}^{\sigma},n_{r}^{\sigma}\}$ is the set of all decision variables. $\Delta_{k}$ is the bound of $d$ in iteration $k$. $g_{i}(\omega_{k})$ represents the left-hand side of standard form inequality constraints (C-4)-(C-7) derived from the original LDP, assuming that the number of robots is less than the available VPs (considering the parking demand of robots during downtime, this setting is reasonable):
$-(n_{h}-1)(n_{v}-1)+N_{o}\leq 0$ (C-4)
[…]
$n_{r}^{\sigma}-\beta(n_{h},n_{v})\cdot(1-{(1-\frac{n_{w}^{\sigma}}{n_{h}+n_{v}})}^{2})\cdot\frac{2\tau_{e}\cdot[n_{h}(n_{v}-1)+n_{v}(n_{h}-1)]}{\tau_{c}}\leq 0\qquad\forall\sigma\in\mathcal{S}$ (C-6)
$-\frac{D\cdot n_{r}^{\sigma}}{\tau_{e}\cdot E[l(n_{h},n_{v},\frac{n_{w}^{\sigma}}{n_{h}+n_{v}})]}+{T}^{\sigma}\leq 0\qquad\forall\sigma\in\mathcal{S}$ (C-7)
[…]
@@
$C_{E}(\omega_{k})=C_{d}(\omega_{k})+\mu\sum_{i}max\{0,g_{i}(\omega_{k})\}$ (C-8)
$C_{EL_{k}}(\omega_{k})=C_{d}(\omega_{k})+\nabla{C_{d}(\omega_{k})}^{T}d_{k}+\mu\sum_{i}max\{0,g_{i}(\omega_{k})+{\nabla g_{i}(\omega_{k})}^{T}\cdot d_{k}\}$ (C-9)
$R_{k}=\frac{C_{E}(\omega_{k})-C_{E}(\omega_{k}+d_{k})}{C_{EL_{k}}(\omega_{k})-C_{EL_{k}}(\omega_{k}+d_{k})}$ (C-10)
where $\mu$ is a large enough constant. Given that the variables in the original problem are integers, we introduced a rounding-up step for $\omega_{k}$ and the bounds in the search process. Additionally, due to the relatively limited solution space, the number of iterations in experiments remained below 50, and the total solving time was in the order of seconds. To ensure the feasibility of the solution for the original problem, the numerical update process includes a rounding-up step.